The company Davis Travel want to know their WACC and to use it for their projects. The following data has to be used to calculate WACC. 20 year bond paying the market rate or Cost of Debt Corporate taxes Preferred stock share par value Preferred stock share market value issuance cost per one preferred share Preferred dividends Common stock share par value Common stock share market value Issuance cost per common stock share is 10% of par value price 19.5 15 % 90 95 92 | 10 Common share dividends last vear | 5 % Constant dividends payment rate (g) on common stock share Capital structure ( 000) Bonds Preferred stock Common stock 2000 1500 3700
Answer #1

Cost of debt = Interest rate(1-tax rate)

=10%(1-0.15)

=10%(0.85)

=8.5%

Cost of preference stock = Dividend/current market price - insurance cost

=7/90-3

=7/87

=8.046%

Cost of equity = D1/(po-insurance cost) + growth rate

D1 = dividend of next year = dividend of current year(1+ growth rate)

=10(1+0.05)

=10(1.05)

=10.5

po= Price of share today = 92

Ke = cost of equity = 10.5/(92 -9.5) + 5%

=10.5/(82.5)+5%

=12.7273%+5%

=17.7273%

Statement showing WACC

Source of capital Amount Weight K WACC = weight*k
Bonds 2000 27.78% 8.50% 2.36%
Prefered stock 1500 20.83% 8.05% 1.68%
common stock 3700 51.39% 17.73% 9.11%
7200 13.15%

Thus WACC = 13.15%